What is the order of operation for this problem. I have tried and keep getting stuck. Here's the problem:
8x + 9y = 31 -5x + y = 26
Here's what i've done so far
8x+5x=31-9y+y=26
13x+31-8y+26 ??????
Thank you

Answer by sudhanshu_kmr(1152)

i hope there are two equations...
8x + 9y = 31 ......(1)
-5x + y = 26 ......(2)
please write each equation in separate line.
multiply equ (1) by 5 and equ (2) by 8 then add both equati....
40x + 45y = 155
-40x + 8y = 208
now, 53 y = 363
y = 363/53 = 6.84
now, u can find the value of x using any one of above euation...